# Conjugacy properties of multivariate unified skew–elliptical distributions

### v2

@@ 2.3 Properties of unified skew–elliptical distributions
$\tilde{{\bm{\Delta}}}_{i\mid j}=\begin{bmatrix}\bar{\bm{\Omega}}_{ij}&{\bm{\Delta}}_{i}\end{bmatrix},\qquad\bar{\tilde{\bm{\Gamma}}}_{i\mid j}=\begin{bmatrix}\bar{\bm{\Omega}}_{jj}&{\bm{\Delta}}_{j}\\ {\bm{\Delta}}_{j}^{\top}&\bar{\bm{\Gamma}}\end{bmatrix},\qquad\tilde{\bm{\tau}}_{i\mid j}=\begin{bmatrix}\bm{\omega}_{j}^{-1}{\bm{\xi}}_{j}\\ \bm{\tau}\end{bmatrix},$ (17)
@@
$\mathbb{P}({\bf{z}}_{j}>\bm{0}\mid{\bf{z}}_{i})=\frac{\mathbb{P}(\bar{{\bf{z}}}_{j}>\bm{0},\bar{{\bf{z}}}_{0}>\bm{0}\mid\bar{{\bf{z}}}_{i}={\bf{z}}_{i})}{\mathbb{P}(\bar{{\bf{z}}}_{0}>\bm{0}\mid\bar{{\bf{z}}}_{i}={\bf{z}}_{i})},\quad\mbox{and}\quad\mathbb{P}({\bf{z}}_{j}>\bm{0})=\frac{\mathbb{P}(\bar{{\bf{z}}}_{j}>\bm{0},\bar{{\bf{z}}}_{0}>\bm{0})}{\mathbb{P}(\bar{{\bf{z}}}_{0}>\bm{0})}=\frac{\mathbb{P}(\bm{\omega}_{j}^{-1}\bar{{\bf{z}}}_{j}>\bm{0},\bar{{\bf{z}}}_{0}>\bm{0})}{\mathbb{P}(\bar{{\bf{z}}}_{0}>\bm{0})}.$ (18)
@@
Leveraging again the closure under linear combinations, marginalization, and conditioning of elliptical distributions [e.g., 35], we have that
@@
Let $\bm{\Gamma}$ be a positive–definite matrix, then ${{\mbox{SUE}}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bm{\Gamma},g^{(m+q)})\stackrel{{\scriptstyle d}}{{=}}{{\mbox{SUE}}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}}\bm{\gamma}^{-1},\bm{\gamma}^{-1}\bm{\tau},\bar{\bm{\Gamma}},g^{(m+q)})$, where $\bar{\bm{\Gamma}}$ is a Pearson correlation matrix defined as $\bar{\bm{\Gamma}}=\bm{\gamma}^{-1}\bm{\Gamma}\bm{\gamma}^{-1}$, with $\bm{\gamma}={\mbox{diag}}(\bm{\Gamma})^{1/2}$.
@@
Let ${\bf{z}}\sim{\mbox{SUE}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bm{\Gamma},g^{(m+q)})$. Then, according to the selection representation in Equation (2), ${\bf{z}}\stackrel{{\scriptstyle d}}{{=}}(\bar{{\bf{z}}}\mid\bar{{\bf{z}}}_{0}>\bm{0})$, and its density function is given by (3). Notice that, since $\bm{\gamma}$ is a diagonal matrix with non–negative entries, the numerator and denominator in (3), can be alternatively re–written as $\mathbb{P}(-\bar{{\bf{z}}}_{0}\leq\bm{0}\mid\bar{{\bf{z}}}={\bf{z}})=\mathbb{P}(-\bm{\gamma}^{-1}\bar{{\bf{z}}}_{0}\leq\bm{0}\mid\bar{{\bf{z}}}={\bf{z}})$ and $\mathbb{P}(-\bar{{\bf{z}}}_{0}\leq\bm{0})=\mathbb{P}(-\bm{\gamma}^{-1}\bar{{\bf{z}}}_{0}\leq\bm{0})$. Therefore, the proof follows directly from (2)–(3) and by the closure under linear combinations and conditioning of elliptical distributions. ∎
@@
with ${\bm{\Delta}}_{i}\in\mathbb{R}^{m\times q_{i}}$, $\bm{\tau}_{i}\in\mathbb{R}^{q_{i}}$, and $\bar{\bm{\Gamma}}_{ij}\in\mathbb{R}^{q_{i}\times q_{j}}$, for every $i,j\in\{1,2\}$, such that $q_{1}+q_{2}=q$. Then, (i) if ${\bm{\Delta}}=\bm{0}$ and $\bm{\tau}=\bm{0}$, it follows that ${\bf{z}}\sim{\mbox{{EC}}}_{m}({\bm{\xi}},\bm{\Omega},g^{(m)})$. Additionally, (ii) if ${\bm{\Delta}}_{i}=\bm{0}$, $\bm{\tau}_{i}=\bm{0}$, $\bar{\bm{\Gamma}}_{ij}=\bm{0}$, for $i$ and $j$ fixed, with $j\neq i$, then ${\bf{z}}\sim{{\mbox{SUE}}}_{m,q_{j}}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}}_{j},\bm{\tau}_{j},\bar{\bm{\Gamma}}_{jj},g^{(m+q_{j})})$. Finally, (iii) $F_{m+q}\{[({\bf{z}}{-}{\bm{\xi}})^{\top}{,}\bm{0}^{\top}]^{\top};{{\mbox{diag}}}(\bm{\Omega},\bar{\bm{\Gamma}}),g^{(m+q)}\}=~{}F_{m}({\bf{z}}{-}{\bm{\xi}};\bm{\Omega},g^{(m)}){\cdot}F_{q}(\bm{0};\bar{\bm{\Gamma}},g^{(q)})$, where ${{\mbox{diag}}}(\bm{\Omega},\bar{\bm{\Gamma}})$ denotes a block–diagonal matrix with blocks $\bm{\Omega}$ and $\bar{\bm{\Gamma}}$, respectively.
@@
To prove (i) note that, due to the invariance of orthant probabilities under centered elliptical distributions [e.g, 35], we have $F_{q}(\bm{0};\bar{\bm{\Gamma}},g^{(q)}_{Q({\bf{z}})})=F_{q}(\bm{0};\bar{\bm{\Gamma}},g^{(q)})$. Hence, including the constraints ${\bm{\Delta}}=\bm{0}$ and $\bm{\tau}=\bm{0}$ in (1), yields
@@
which coincides with the density of the elliptical distribution $\mbox{{EC}}_{m}({\bm{\xi}},\bm{\Omega},g^{(m)})$. Such a result allows us to prove also (iii). In particular, since ${\bf{z}}\sim{\mbox{{EC}}}_{m}({\bm{\xi}},\bm{\Omega},g^{(m)})$ when both ${\bm{\Delta}}=\bm{0}$ and $\bm{\tau}=\bm{0}$, then $\mathcal{P}({\bf{z}})=F_{m}({\bf{z}}-{\bm{\xi}};\bm{\Omega},g^{(m)})$. Conversely, by including the constraints ${\bm{\Delta}}=\bm{0}$ and $\bm{\tau}=\bm{0}$ within the general expression for the sue cumulative distribution function in (4) yields $\mathcal{P}({\bf{z}})=F_{m+q}\{[({\bf{z}}-{\bm{\xi}})^{\top},\bm{0}^{\top}]^{\top};{{\mbox{diag}}}(\bm{\Omega},\bar{\bm{\Gamma}}),g^{(m+q)}\}/F_{q}(\bm{0};\bar{\bm{\Gamma}},g^{(q)})$. Therefore $\mathcal{P}({\bf{z}})=F_{m}({\bf{z}}-{\bm{\xi}};\bm{\Omega},g^{(m)})=F_{m+q}\{[({\bf{z}}-{\bm{\xi}})^{\top},\bm{0}^{\top}]^{\top};{{\mbox{diag}}}(\bm{\Omega},\bar{\bm{\Gamma}}),g^{(m+q)}\}/F_{q}(\bm{0};\bar{\bm{\Gamma}},g^{(q)})$ which implies the result stated in point (iii).
@@
In the particular case of a sun distribution, it can be shown that the restrictions $\bm{\tau}=\bm{0}$ or $\bm{\tau}_{i}=\bm{0}$ are unnecessary in Lemma 5 since there are results analogous to (i)—(iii) that follow directly by the specific properties of Gaussian cumulative distribution functions. In particular, adapting the above proof to the sun sub–family, it can be easily shown, for example, that if ${\bm{\Delta}}_{i}=\bm{0}$ and $\bar{\bm{\Gamma}}_{ij}=\bar{\bm{\Gamma}}^{\top}_{ji}=\bm{0}$, $j\neq i$, then ${\bf{z}}\sim{{\mbox{SUN}}}_{m,q_{j}}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}}_{j},\bm{\tau}_{j},\bar{\bm{\Gamma}}_{jj})$, even if $\bm{\tau}_{i}\neq\bm{0}$.
@@
3 Conjugacy properties of multivariate unified skew–elliptical (sue) distributions
@@
Sections 3.13.3 present the novel results on the conjugacy properties of the sue family under a broad class of statistical models for fully–observed, censored or dichotomized realizations from elliptical or skew–elliptical variables. As anticipated in Section 1, the technical derivation of these results is based on specifying a general joint sue distribution for the parameters ${\bm{\beta}}$ and the noise vector ${\bm{\varepsilon}}$ underlying the response ${\bf{y}}$. This allows to leverage the closure properties in Section 2.3, to obtain closed–form sue priors $\mathcal{p}({\bm{\beta}})$ and meaningful likelihoods $\mathcal{p}({\bf{y}}\mid{\bm{\beta}})$ whose combination, under the standard Bayes rule, yields posterior distributions $\mathcal{p}({\bm{\beta}}\mid{\bf{y}})\propto\mathcal{p}({\bm{\beta}})\mathcal{p}({\bf{y}}\mid{\bm{\beta}})$ that still belong to the sue class.
@@
The above technical focus on the joint distribution $\mathcal{p}({\bm{\beta}},{\bf{y}})$ for the data ${\bf{y}}$ and the parameters ${\bm{\beta}}$ is motivated by the fact that the results in Section 3.13.3 crucially clarify that not all models arising from elliptical or skew–elliptical noise vectors admit conjugate sue priors. For this property to hold generally within the sue family, it is necessary to consider a form of dependence between ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ due to the specific properties of the density generator. Notice that such a dependence is often weak. In particular, it allows to account for meaningful priors and models having ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ uncorrelated, while reducing to full independence under the density generators of the multivariate Gaussians and unified skew–normals. Nonetheless, such a weak dependence combined with a technical focus on $\mathcal{p}({\bm{\beta}},{\bf{y}})$ allows for a more comprehensive investigation of sue conjugacy properties that would not be as immediate to prove theoretically under a direct specification of the prior $\mathcal{p}({\bm{\beta}})$ and the likelihood $\mathcal{p}({\bf{y}}\mid{\bm{\beta}})$.
@@
As a simple illustrative example that clarifies the above arguments, consider a univariate setting with $\mbox{Cauchy}(0,1)$ prior for $\beta$ and a $\mbox{Cauchy}(\beta,1)$ likelihood for $(y\mid\beta)$. By application of the Bayes rule, we obtain $\mathcal{p}(\beta\mid y)\propto\mathcal{p}(\beta)\mathcal{p}(y\mid\beta)$ where $\mathcal{p}(\beta)\mathcal{p}(y\mid\beta)=1/[\pi^{2}(1+\beta^{2})(1+(y-\beta)^{2})]$, which is not proportional to the kernel of a Cauchy density. Clearly, in this simple example and in general situations where conjugacy lacks, Bayesian inference can still proceed via routinely–implemented mcmc methods or deterministic approximations of the target posterior. Nonetheless, as clarified in Sections 3.13.3, sue conjugacy can be still achieved under certain likelihoods induced by elliptical or skew–elliptical error terms (including instances of potential practical interest), thereby facilitating posterior inference.

### v3

@@ 2.3 Properties of unified skew-elliptical distributions
$\tilde{{\bm{\Delta}}}_{i\mid j}=\begin{bmatrix}\bar{\bm{\Omega}}_{ij}&{\bm{\Delta}}_{i}\end{bmatrix},\qquad\tilde{\bm{\tau}}_{i\mid j}=\begin{bmatrix}\bm{\omega}_{j}^{-1}{\bm{\xi}}_{j}\\ \bm{\tau}\end{bmatrix},\qquad\bar{\tilde{\bm{\Gamma}}}_{i\mid j}=\begin{bmatrix}\bar{\bm{\Omega}}_{jj}&{\bm{\Delta}}_{j}\\ {\bm{\Delta}}_{j}^{\top}&\bar{\bm{\Gamma}}\end{bmatrix},$ (17)
@@
$\mathbb{P}({\bf{z}}_{j}>\bm{0}\mid{\bf{z}}_{i})=\frac{\mathbb{P}(\bar{{\bf{z}}}_{j}>\bm{0},\bar{{\bf{z}}}_{0}>\bm{0}\mid\bar{{\bf{z}}}_{i}={\bf{z}}_{i})}{\mathbb{P}(\bar{{\bf{z}}}_{0}>\bm{0}\mid\bar{{\bf{z}}}_{i}={\bf{z}}_{i})},\qquad\mathbb{P}({\bf{z}}_{j}>\bm{0})=\frac{\mathbb{P}(\bar{{\bf{z}}}_{j}>\bm{0},\bar{{\bf{z}}}_{0}>\bm{0})}{\mathbb{P}(\bar{{\bf{z}}}_{0}>\bm{0})}=\frac{\mathbb{P}(\bm{\omega}_{j}^{-1}\bar{{\bf{z}}}_{j}>\bm{0},\bar{{\bf{z}}}_{0}>\bm{0})}{\mathbb{P}(\bar{{\bf{z}}}_{0}>\bm{0})}.$ (18)
@@
Leveraging again the closure under linear combinations, marginalization and conditioning of elliptical distributions [e.g., 35], we have that
@@
Let $\bm{\Gamma}$ be a positive-definite matrix, then ${{\mbox{SUE}}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bm{\Gamma},g^{(m+q)})\stackrel{{\scriptstyle d}}{{=}}{{\mbox{SUE}}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}}\bm{\gamma}^{-1},\bm{\gamma}^{-1}\bm{\tau},\bar{\bm{\Gamma}},g^{(m+q)})$, where $\bar{\bm{\Gamma}}$ is a Pearson correlation matrix defined as $\bar{\bm{\Gamma}}=\bm{\gamma}^{-1}\bm{\Gamma}\bm{\gamma}^{-1}$, with $\bm{\gamma}={\mbox{diag}}(\bm{\Gamma})^{1/2}$.
@@
Let ${\bf{z}}\sim{\mbox{SUE}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bm{\Gamma},g^{(m+q)})$. Then, according to the selection representation in (2), ${\bf{z}}\stackrel{{\scriptstyle d}}{{=}}(\bar{{\bf{z}}}\mid\bar{{\bf{z}}}_{0}>\bm{0})$, and its density function is given by (3). Notice that, since $\bm{\gamma}$ is a diagonal matrix with non-negative entries, then the numerator and denominator in (3) can be alternatively re-written as $\mathbb{P}(-\bar{{\bf{z}}}_{0}\leq\bm{0}\mid\bar{{\bf{z}}}={\bf{z}})=\mathbb{P}(-\bm{\gamma}^{-1}\bar{{\bf{z}}}_{0}\leq\bm{0}\mid\bar{{\bf{z}}}={\bf{z}})$ and $\mathbb{P}(-\bar{{\bf{z}}}_{0}\leq\bm{0})=\mathbb{P}(-\bm{\gamma}^{-1}\bar{{\bf{z}}}_{0}\leq\bm{0})$. Therefore, the proof follows directly from (2)–(3) and by the closure under linear combinations and conditioning of elliptical distributions. ∎
@@
with ${\bm{\Delta}}_{i}\in\mathbb{R}^{m\times q_{i}}$, $\bm{\tau}_{i}\in\mathbb{R}^{q_{i}}$, and $\bar{\bm{\Gamma}}_{ij}\in\mathbb{R}^{q_{i}\times q_{j}}$, for every $i,j\in\{1,2\}$, such that $q_{1}+q_{2}=q$. Then, (i) if ${\bm{\Delta}}=\bm{0}$ and $\bm{\tau}=\bm{0}$, it follows that ${\bf{z}}\sim{\mbox{{EC}}}_{m}({\bm{\xi}},\bm{\Omega},g^{(m)})$. Additionally, (ii) if ${\bm{\Delta}}_{i}=\bm{0}$, $\bm{\tau}_{i}=\bm{0}$, $\bar{\bm{\Gamma}}_{ij}=\bm{0}$, for $i$ and $j$ fixed, with $j\neq i$, then ${\bf{z}}\sim{{\mbox{SUE}}}_{m,q_{j}}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}}_{j},\bm{\tau}_{j},\bar{\bm{\Gamma}}_{jj},g^{(m+q_{j})})$. Finally, (iii) $F_{m+q}\{[({\bf{z}}-{\bm{\xi}})^{\top},\bm{0}^{\top}]^{\top};{{\mbox{diag}}}(\bm{\Omega},\bar{\bm{\Gamma}}),g^{(m+q)}\}=F_{m}({\bf{z}}-{\bm{\xi}};\bm{\Omega},g^{(m)})\cdot F_{q}(\bm{0};\bar{\bm{\Gamma}},g^{(q)})$, where ${{\mbox{diag}}}(\bm{\Omega},\bar{\bm{\Gamma}})$ denotes a block-diagonal matrix with blocks $\bm{\Omega}$ and $\bar{\bm{\Gamma}}$, respectively.
@@
To prove (i) note that, due to the invariance of orthant probabilities under centered elliptical distributions [e.g., 35], we have $F_{q}(\bm{0};\bar{\bm{\Gamma}},g^{(q)}_{Q({\bf{z}})})=F_{q}(\bm{0};\bar{\bm{\Gamma}},g^{(q)})$. Hence, including the constraints ${\bm{\Delta}}=\bm{0}$ and $\bm{\tau}=\bm{0}$ in (1), yields
@@
which coincides with the density of the elliptical distribution $\mbox{{EC}}_{m}({\bm{\xi}},\bm{\Omega},g^{(m)})$. This result allows us to prove also (iii). In particular, since ${\bf{z}}\sim{\mbox{{EC}}}_{m}({\bm{\xi}},\bm{\Omega},g^{(m)})$ when both ${\bm{\Delta}}=\bm{0}$ and $\bm{\tau}=\bm{0}$, then $\mathcal{P}({\bf{z}})=F_{m}({\bf{z}}-{\bm{\xi}};\bm{\Omega},g^{(m)})$. Conversely, by including the constraints ${\bm{\Delta}}=\bm{0}$ and $\bm{\tau}=\bm{0}$ within the general expression for the SUE cumulative distribution function in (4) yields $\mathcal{P}({\bf{z}})=F_{m+q}\{[({\bf{z}}-{\bm{\xi}})^{\top},\bm{0}^{\top}]^{\top};{{\mbox{diag}}}(\bm{\Omega},\bar{\bm{\Gamma}}),g^{(m+q)}\}/F_{q}(\bm{0};\bar{\bm{\Gamma}},g^{(q)})$. Therefore $\mathcal{P}({\bf{z}})=F_{m}({\bf{z}}-{\bm{\xi}};\bm{\Omega},g^{(m)})=F_{m+q}\{[({\bf{z}}-{\bm{\xi}})^{\top},\bm{0}^{\top}]^{\top};{{\mbox{diag}}}(\bm{\Omega},\bar{\bm{\Gamma}}),g^{(m+q)}\}/F_{q}(\bm{0};\bar{\bm{\Gamma}},g^{(q)})$ which implies the result stated in point (iii).
@@
In the particular case of a SUN distribution, it can be shown that the restrictions $\bm{\tau}=\bm{0}$ or $\bm{\tau}_{i}=\bm{0}$ are unnecessary in Lemma 5 since there are results analogous to (i)–(iii) that follow directly by the specific properties of Gaussian cumulative distribution functions. In particular, adapting the above proof to the SUN sub-family, it can be easily shown, for example, that if ${\bm{\Delta}}_{i}=\bm{0}$ and $\bar{\bm{\Gamma}}_{ij}=\bar{\bm{\Gamma}}^{\top}_{ji}=\bm{0}$, $j\neq i$, then ${\bf{z}}\sim{{\mbox{SUN}}}_{m,q_{j}}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}}_{j},\bm{\tau}_{j},\bar{\bm{\Gamma}}_{jj})$, even if $\bm{\tau}_{i}\neq\bm{0}$.
@@
3 Conjugacy properties of multivariate unified skew-elliptical (SUE) distributions
@@
Sections 3.13.3 present the new results on the SUE conjugacy properties under a broad class of regression models for fully-observed, censored or dichotomized realizations from elliptical or skew-elliptical variables. As anticipated in Section 1, the technical derivation of these results is based on specifying a general joint SUE distribution for the parameters ${\bm{\beta}}$ and the noise vector ${\bm{\varepsilon}}$ underlying the response ${\bf{y}}$. This allows to leverage the closure properties in Section 2.3, to obtain closed-form SUE priors $\mathcal{p}({\bm{\beta}})$ and meaningful likelihoods $\mathcal{p}({\bf{y}}\mid{\bm{\beta}})$ whose combination, under the standard Bayes rule, yields posterior distributions $\mathcal{p}({\bm{\beta}}\mid{\bf{y}})\propto\mathcal{p}({\bm{\beta}})\mathcal{p}({\bf{y}}\mid{\bm{\beta}})$ that still belong to the SUE class.
@@
The above technical focus on the joint distribution $\mathcal{p}({\bm{\beta}},{\bf{y}})$ for the data ${\bf{y}}$ and the parameters ${\bm{\beta}}$ is motivated by the fact that the results in Section 3.13.3 crucially clarify that not all the models arising from elliptical or skew-elliptical noise vectors admit conjugate SUE priors. For this property to hold generally within the SUE family, it is necessary to consider a form of dependence between ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ due to the specific properties of the density generator. Notice that such a dependence is often weak. In particular, it allows to account for meaningful priors and models having ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ uncorrelated, while reducing to full independence under the density generators of the multivariate Gaussians and unified skew-normals. Nonetheless, such a weak dependence combined with a technical focus on $\mathcal{p}({\bm{\beta}},{\bf{y}})$ allows for a more comprehensive investigation of SUE conjugacy properties that would not be as immediate to prove theoretically under a direct specification of the prior $\mathcal{p}({\bm{\beta}})$ and the likelihood $\mathcal{p}({\bf{y}}\mid{\bm{\beta}})$.
@@
As a simple illustrative example that clarifies the above arguments, consider a univariate setting with $\mbox{Cauchy}(0,1)$ prior for $\beta$ and a $\mbox{Cauchy}(\beta,1)$ likelihood for $(y\mid\beta)$. By application of the Bayes rule, we obtain $\mathcal{p}(\beta\mid y)\propto\mathcal{p}(\beta)\mathcal{p}(y\mid\beta)$ where $\mathcal{p}(\beta)\mathcal{p}(y\mid\beta)=1/[\pi^{2}(1+\beta^{2})(1+(y-\beta)^{2})]$ is not proportional to the kernel of a Cauchy density. Clearly, in this simple example and in general situations where conjugacy lacks, Bayesian inference can still proceed via routinely-implemented MCMC methods or deterministic approximations of the target posterior. Nonetheless, as clarified in Sections 3.13.3, SUE conjugacy can be still achieved under certain likelihoods induced by elliptical or skew-elliptical error terms (including instances of potential practical interest), thereby facilitating posterior inference.